ICTS Project Update
Ricoh Invoice Automation and Telephone Usage Billing
ICTS has been working with the vendor to install an API that will allow seamless integration between Talend (MIT’s Middleware system) and Technology One (MIT’s finance system). Once the API is fully installed and operational, MIT will engage another vendor to develop the code that sits within Talend to automate the entire process. The invoice sorting and report mechanisms are completed – this work just needs to be transferred into Talend.
Business Intelligence Investigation
During the month of December, the BI business case was signed off and approved. ICTS worked with the vendor in installing the technical platform and configuring all the servers. The vendor has also done significant worn with Finance to develop an online EFT report as a pilot project. Once this report is signed off, ICTS will work with the PMO in setting up a BI working group and push the application to the wider MIT community. Once this report is signed off, ICTS will work with the PMO in setting up a BI working group and push the application to a wider MIT community.
Roll-out of Skype for Business to end-users now complete apart from MIT Enterprise. Remaining tasks before switching off Mitel are resolving outstanding 488 issue, replacing the fax capability, migrate analogue phones to AudioCode and replacement of Contact Centre.
Migration of WAN & Internet Services to 2degrees
Middlemore successfully cut-over, project will now be closed following project review.
ICTS is happy to report that the Heat upgrade took place successfully. ICTS is however planning to review the tool holistically in 2017 to ensure its compliance to MIT requirements.
MIT Enterprise Network Upgrade
Progressing according to project plan.
Network Management & Monitoring – SOW to be finalised, hardware delivered.
2016 CAPEX Device replacement
2016 replacement plan has been produced from our Service Desk asset database, with replacing oldest equipment first and sent to Faculty and Service Centre Managers. All equipment is onsite and the Service Delivery Team is working through a priority list coupled with availability. Classrooms are the first priority so equipment is ready for teaching mid January.
BA work completed awaiting FoB agreement on migration dates for remaining groups.
Backup & Archiving
Implementation of Panzura underway. Will take several ‘months’ to migrate data.
Replacement Contact Centre Software
Evaluation team have short-listed two products. Preferred product selected and recommendation will be actioned in January.
- Received draft information from Facilities Management in regards to the failed test of MITM Data Centre Bypass Switch on the 4th October. FM are waiting on further information from consultants before any retest.
- Dell KACE environment decommissioned
- Urgent request to apply two latest hotfixes to ebs_test and production of EBS completed 5th December
- HEAT upgraded to version 2016.1 on 6th December
- Continued with network switch replacement program throughout December.
- eMIT switched off 17th December
- Continued with Patching program including Skype environment
- Retired old print servers on the 18th December.
*Firewall stats for December are not available at this stage*
ICTS Service Level Agreement Performance
Enterprise Services Availability
All Enterprises Services were at 99.5% or higher availability for December.
Services availability is measured on total uptime (excluding planned scheduled outage windows) minus any unscheduled outages.
The Availability measurement for services listed is 7.30am to 6.00pm, 5 days a week apart from Voyager and this measurement is based on 7.30 to 9.00pm, 7 days a week. Core Infrastructure Services (Compute and Network) are available 24 hours, 7 days a week.
Major Infrastructure Outages/Incidents during December 2016
There were no service outage during SLA hours in December.
Issues during December:
- Issues between DMZ switch and Untrusted Fortigate Firewall in NB on the 5th December. Problem fixed by replacing 5M – 10GB twinax cable.
- Unscheduled outage of Voyager – the main library system. Voyager is fully managed by the developers, ExLibris and it appears the planned work for Thursday evening was actioned a day early. Issue managed by Library team.
- Network switch problem at Maritime on 20th December where main switch appeared to lose stack configuration. Reloaded configuration.
- Broken switch in NP block on 20th December impacted the wireless service. Replaced switch and services restored.
Average size of Email boxes/storage
Incident and Service Request Reporting
Actual Incident Summary
Actual Service Request Summary
Service Level Targets